# quality of service 5G fifth generation

### v2

@@ IV Methodology
IV-A Exhaustive Search Algorithm
The ESA is a fundamental approach used to find the global optimal solution to resource allocation problems by evaluating all possible configurations and selecting the best one based on a predefined objective function [33]. The ESA is a valuable benchmark solution that provides a ground truth against comparing results from other methods. Its advantages include simplicity of understanding and implementation and the assurance of yielding an optimal solution within a finite time frame. However, its guarantee of optimality comes at the cost of high computational effort, making it suitable primarily for small or well-defined problems where the solution space can be thoroughly explored.
Nonetheless, ESA’s main limitation lies in its computational complexity, especially when addressing resource allocation problems with high dimensions and complex constraints. The computational cost increases exponentially as the problem size grows, making ESA impractical for large-scale resource allocation tasks [34]. Despite this limitation, ESA is still widely used as a benchmark for comparing different optimization algorithms and as a reference for evaluating new optimization techniques.
To evaluate the performance, accuracy, and reliability of the SS-VAE method, we will compare its results with those obtained by ESA. We will use the ESA to obtain a subset of solutions for the objective variables ($\alpha$, $\beta$, $p$) in the resource allocation optimization problem with network slicing. This labeled dataset will be used to train the supervised learning part of our SS-VAE method. The detailed steps of ESA are described in Algorithm 1.
@@ IV-A Exhaustive Search Algorithm
IV-B Deep Reinforcement learning Algorithm
As discussed in Section II, DRL algorithms effectively address resource allocation, proving their practicality in wireless networks and enhancing key metrics [35]. In this study, we use DRL, where an agent allocates resources based on rewards linked to the objective in Problem 10. Hence, we model the problem as a Markov Decision Process (MDP) and use a DRL algorithm as a benchmark for evaluating our SS-VAE method. In the resource allocation problem (Equation 10), an O-RAN orchestrator is responsible for determining the PRB allocation decisions and identifying the corresponding RU for each UE within each slice. This problem can be represented as a MDP [35, 36, 37, 38].
@@ IV-B Deep Reinforcement learning Algorithm
    $\mathfrak{R(s,a)}=\Theta_{r}\;R_{u,s}+\Theta_{const.}\;C_{u,b,m,s}+\Theta_{bias},$ (22)
@@
The corresponding value function $V(\pi):\mathfrak{S}\rightarrow\mathfrak{R}$ is defined as the cumulative discounted reward achieved by following the actions prescribed by the policy $\pi$ when starting from a specific initial state. Exploring the dynamics of the environment, especially when accounting for variable channel conditions and calculating transition probabilities, is a highly complex task. Determining the value function in such scenarios presents a significant challenge. Two primary strategies are employed to address MDPs: model-based and model-free approaches. In uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
Our resource allocation problem involves managing high-dimensional and complex state and action spaces due to continuous variables like transmission power and numerous UEs and PRB resources. Traditional Q-learning methods struggle in this context because of inefficient learning with rarely encountered state-action pairs [35, 38]. To solve the problem of finding parameters such as transmission power ($p$), $\alpha$, and $\beta$, we use the Deep Q-Network (DQN) approach, which handles complex state-action spaces by approximating the $Q$ function with a deep neural network. DQN benefits from experience replay and target networks, enhancing learning stability and efficiency [38]. Experience replay stores transition in memory ($\mathbb{M}$) and samples them to train the neural network ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$), reducing correlation in the data and improving gradient estimation. A target network stabilizes learning by updating the $Q$-network’s parameters with independent samples [38].
Our DQN training employs the epsilon-greedy policy for action selection. This means that the agent chooses the action with the highest $Q$-value with a probability of $1-\epsilon$, and with a probability of $\epsilon$, it explores and selects a random action. The policy function is represented as:
@@
It is worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
IV-C Proposed Method (SS-VAE)
Traditional resource allocation methods in wireless communication systems, such as heuristic and meta-heuristic algorithms, often rely on predefined rules or optimization algorithms that do not adapt well to changing network conditions. In recent times, DL techniques have been used to improve the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems [39, 40, 41, 42]. These approaches are practical in learning complex, non-linear relationships between input features (e.g., channel matrices) and output decisions (e.g., power and resource allocation).
The distinctive feature of DL-based resource allocation is its capacity to manage wireless communication systems’ dynamic and complex nature effectively. DL-based approaches can adjust to changes in network conditions, user demands, and interference patterns, providing more flexible and efficient resource management than traditional methods. Traditional methods often face challenges in real-time adaptation due to predefined rules, complexity in optimization, and a lack of real-time learning. Our SS-VAE method inputs the channel gain matrix into a DNN. This network learns optimal resource allocation decisions that satisfy network constraints and maximize the weighted aggregated data rate. We use a flattened vector, with dimensions $UB$, derived from the channel gain matrix as the input for the machine learning model.
Our main contribution involves integrating a VAE with contrastive loss to solve the resource allocation problem. By adding contrastive loss into the VAE‘s loss function, we enhance the encoder’s ability to learn more refined representations, significantly improving the accuracy of resource allocation decisions. The SS-VAE method allows the encoder layers of the VAE to be explicitly utilized for the resource allocation task.
In our problem, the connection between the xApps that handle power transmissions and PRB allocation emerges from their shared foundation within the VAE framework. The VAE operates as a shared space where both xApps come together, enabling a collaborative exploration and consideration of the relevant features for both power transmission and PRB allocations.
The xApps access a common latent representation of the channel gain matrices through this shared space. This shared understanding facilitates a cohesive analysis where the relationships between power transmissions and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the key features that impact both domains. This integration of information helps make more comprehensive and informed decisions about resource allocation, as it ensures that both xApps use a unified understanding of the underlying data to optimize their respective objectives. This coordination allows for a synchronized approach to resource allocation within the O-RAN architecture.
IV-C1 Variational Autoencoder
The VAE is a type of DNN generative model used in unsupervised learning tasks. It is designed to generate new data samples that resemble a provided set of training data. One notable advantage of using a VAE is its ability to adeptly capture the intrinsic distribution of the channel gain matrix (input data), which is valuable when addressing the resource allocation problem [43].
In the VAE part of our method, the input data (flattened channel gain matrix) is represented by $x$, while a lower-dimensional data representation (latent space) is denoted as $z$. The generative model based on the latent variable is represented as $p_{\theta}(x)$. The main objective of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$. This likelihood can be expressed mathematically as:
@@ IV-C1 Variational Autoencoder
where $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable based on the input data, the term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable. The loss function (Equation 23) can be optimized by adjusting the parameters of the generative model ($\theta$) and the encoder network represented ($\phi$). This optimization process involves iteratively updating these parameters and computing the gradient of the objective with respect to them [43]. The architecture of the VAE for addressing the resource allocation problem with network slicing is illustrated Figure 3. The VAE comprises an encoder and a decoder. The encoder compresses the channel gain matrix into a lower-dimensional latent code while the decoder reconstructs the channel gain matrix from this code. During training, the VAE minimizes reconstruction loss, improving the accuracy of predicted transmission power.
@@
Several key advantages can be observed when using the VAE for resource allocation. The model can represent complex distributions, capture the non-linear relationship between channel gains and resource allocation decisions, and provide a compressed latent space helpful for visualization, clustering, and anomaly detection tasks. Moreover, resource allocation problems are inherently non-convex and challenging for traditional optimization methods, where the VAEs can propose a more versatile and practical solution to addressing these complexities.
The dataset used to train the SS-VAE model consists of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples. These samples are represented by pairs of channel matrices and their corresponding power allocation vectors, denoted as $H_{i}$ and $\Gamma_{i}$ for labeled samples (where $i$ ranges from 1 to $Q_{L}$), and $H_{j}$ for unlabeled samples (where $j$ ranges from 1 to $Q_{U}$). The labeled dataset results from the ESA applied to the same training dataset. Specifically, we have used power transmission and allocation decisions ($p,\alpha,\beta$) obtained by ESA as labels for training our SS-VAE model.
To train the supervised learning component of our algorithm, we employ the ADAM optimizer [44]. It facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples. The loss function for the regression problem with ADAM optimization is formulated using the mean squared error (MSE) loss function, defined as follows:
@@
The encoder part of the VAE is used to estimate the transmission power and predict the allocation indicators ($\alpha_{u,b,s}$ and $\beta_{u,b,m,s}$). Our model takes a flattened vector of dimension $UB$ as input, obtained from the channel gain matrix.
IV-C2 Contrastive Loss
Contrastive loss is a widely-used loss function in DL for various representation learning tasks, including image or text classification. Its fundamental aim is to construct a representation of data points that provokes similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17]. Mathematically, the contrastive loss can be expressed as follows:
@@ IV-C2 Contrastive Loss
where $N$ is the number of data points, $\mathbf{x}_{i}$ is the representation of the $i^{th}$ data point, $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ are the representations of the similar and dissimilar data points respectively, $y_{i}$ is a binary label indicating whether $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$ are similar or dissimilar, and $d(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$ is the distance function between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$. The parameter $m$ is a margin that separates the similar and dissimilar representations.
To apply contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices. This can be achieved using a channel’s statistical model that reflects the relevant properties of the communication environment. By introducing randomness through a Rayleigh fading model, we can produce channel gain matrices with varying coefficients in magnitude and phase, thus creating either similar or dissimilar matrices. To generate a channel gain matrix using this approach, we create a matrix of complex Gaussian random variables with zero mean, where the variance is influenced by the distance between the transmitter and receiver while also accounting for interference as noise [45, 46]. The variance can be adjusted to control the level of correlation between the channels, thereby generating similar or dissimilar channel gain matrices as needed. We can train our model with contrastive loss function once we have created similar and dissimilar channel gain matrices. The loss function can be formulated as:
@@
where $\theta$ is the set of parameters in the VAE model, $\tau$ is a temperature hyperparameter, and $L(\theta)$ is the contrastive loss function. $\underline{H_{i}}$ and $\bar{H_{i}}$ are similar matrices to the main channel gain matrix ($H_{i}$) and are used in training the SS-VAE model [47]. By minimizing this loss function, we can find a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices; therefore, it is used to solve the resource allocation problem.
The connection between the VAE and the additional contrastive loss is essential for an optimal solution. The initial step establishes a strong base by using supervised learning to acquire a good representation of the resource allocation problem. The subsequent step further improves these representations, enhancing the model’s capacity to generalize and effectively handle new, unforeseen scenarios.
IV-C3 Training the SS-VAE model
The training phase of SS-VAE method consists of two parts. First, supervised learning involves training the VAE and using the VAE’s Encoder to predict power transmission. The second part is unsupervised learning, which uses contrastive loss to improve the generalization and robustness of our method.
The latent space dimension represents the size of the hidden representation learned by the VAE [48]. A larger latent space dimension can lead to a more expressive model but also increases the risk of overfitting. The learning rate determines the speed at which the model learns from the data. The batch size determines the number of samples used for each iteration of the training process. The number of epochs determines the number of times the model will be trained on the data. $L1$ and $L2$ regularization to prevent overfitting by adding a penalty term to the loss function of supervised learning [49]. The model is developed using the PyTorch [50] software framework. The dataset’s validation and test sizes are $20\%$, allowing for a more thorough evaluation of the model’s performance. Hyperparameter tuning yields the value of hyperparameters, which can be found in Table IV.
IV-C4 Performance Metrics
@@ V SIMULATION RESULTS
V-A Initial points and performance of training the model
This section presents the numerical results for the resource allocation problem to evaluate the performance of the SS-VAE model compared to ESA and DQN. In our training and testing scenarios, each RU in the O-RAN deployment is equipped with 25 PRBs and operates with a single antenna. We consider a network with two slices, eMBB and URLLC services. These UEs are randomly distributed within a single-cell environment, mostly 200 to 500 meters from the RU. The initial values and parameters used to run and simulate SS-VAE and the benchmark algorithms (ESA and DQN) are provided in Table IV, Table IV, and Table IV.
TABLE II: Resource Allocation Simulation Parameters
@@ V-A Initial points and performance of training the model
Maximum fronthaul capacity 46 bps
Maximum data rate for eMBB 20 bps
Maximum data rate for URLLC 2 bps
@@
TABLE III: DQN Training Parameters
@@
$\gamma$, $\epsilon_{finalvalue}$, $\epsilon_{decayfactor}$ 0.9, 0.1, 0.9995
@@
TABLE IV: SS-VAE Model Training Parameters

### v3

@@ 5 Methodology
The robustness and efficacy of the SS-VAE framework will be rigorously evaluated using multiple regression-based metrics suitable for assessing the accuracy of the generated resource allocation parameters. Its performance will be benchmarked against both the globally optimal solutions derived from ESA and deployment-ready, state-of-the-art methods such as DRL agents. This comprehensive validation aims to demonstrate the reliability of SS-VAE across various operational scenarios, particularly highlighting its capability to approach optimal performance without the prohibitive computational cost associated with ESA. Figure 3 illustrates our proposed methodology, which encompasses the training pipeline for SS-VAE and a comparative evaluation against benchmark algorithms.
5.1 Exhaustive Search Algorithm (ESA)
We employ ESA method as a basis of comparison for our algorithm. The ESA is a fundamental method to find the global optimal solution for resource allocation problems by exhaustively evaluating all possible configurations based on a predefined objective function [40]. It serves as a benchmark by providing a ground truth for comparing other methods. While it is simple to understand and guarantees optimality within a finite time, its high computational cost limits its use to small or well-defined problems [41]. Nevertheless, it remains a widely used benchmark for evaluating and comparing optimization algorithms.
We use the ESA to create a labeled dataset by obtaining a subset of optimal solutions for the objective variables (i.e., $\alpha$, $\beta$, $p$) in the resource allocation optimization problem. This dataset is subsequently used to train the supervised learning component of our SS-VAE method. The detailed steps of the ESA are described in Algorithm 1.
@@ 5.1 Exhaustive Search Algorithm (ESA)
5.2 Deep Reinforcement learning Algorithm
To evaluate the proposed SS-VAE methodology, we benchmark its performance against a DRL model, a paradigm increasingly recognized for its efficacy in navigating complex resource allocation tasks within contemporary wireless networks [42]. The resource allocation challenge, as formulated in Problem 10, is cast as a Markov Decision Process (MDP). This formalization allows the O-RAN orchestrator, responsible for PRB assignments and RU associations per slice, to operate as a learning agent [42, 43, 44, 45]. The MDP is characterized by:
@@ 5.2 Deep Reinforcement learning Algorithm
    $\mathfrak{R(s,a)}=\Theta_{r}\;R_{u_{s}}+\Theta_{const.}\;C_{u_{s},m}^{b}+\Theta_{bias},$ (22)
@@
The inherent complexity of wireless resource allocation, characterized by high-dimensional state and action spaces and continuous variables (e.g., transmission power), renders traditional Q-learning approaches inefficient due to the curse of dimensionality and issues with generalizing from sparsely visited state-action pairs [42, 45]. We employ a DQN framework to surmount these challenges. DQN leverages a DNN to approximate the action-value function (Q-function), enabling it to manage large-scale problems effectively. Crucially, DQN incorporates mechanisms such as experience replay—storing transitions ($\mathfrak{S}_{t},\mathfrak{A}_{t},\mathfrak{R}_{t},\mathfrak{S}_{t+1}$) in a replay memory ($\mathbb{M}$) to break data correlations—and target networks to stabilize the learning process and improve convergence.
To ensure a robust exploration-exploitation balance during training, the DQN agent utilizes an epsilon-greedy strategy for action selection:
@@
This allows the agent to exploit its current knowledge by selecting the action with the highest estimated Q-value with probability ($1-\epsilon$) or to explore the action space randomly with probability ($\epsilon$). The comprehensive training algorithm for the DQN is detailed, and our specific implementation for this problem is encapsulated in Algorithm 3 [45].
5.3 Semi-Supervised VAE (SS-VAE)
To address the resource allocation problem, we introduce our SS-VAE method in this subsection. The model is designed to process a flattened channel gain matrix (dimension $U\times B$) through a DNN to find optimal resource allocation strategies. The central contribution of our method is the integration of a VAE with a contrastive loss, which improves the learning of refined representations..
In our framework, xApps for power transmission and PRB allocation coordinate through a shared VAE foundation. This VAE acts as a common space, allowing xApps to collaboratively consider relevant features by accessing a shared latent representation of channel gain matrices. This shared understanding facilitates a cohesive analysis and simultaneous consideration of key features impacting both domains. This leads to more comprehensive and informed resource allocation decisions and enables synchronized O-RAN resource management.
5.3.1 Variational Autoencoder
The VAE, a DNN generative model for unsupervised learning, excels at capturing the intrinsic distribution of input data like the channel gain matrix, which is valuable for the resource allocation problem [46]. It maps input data $x$ (flattened channel gain matrix) to a lower-dimensional latent space $z$, with $p_{\theta}(x)$ as the generative model. The VAE aims to maximize data likelihood given model parameters $\theta$, as expressed mathematically by:
@@ 5.3.1 Variational Autoencoder
where $q_{\phi}(z|x)$ is the encoder to model the posterior distribution of the latent variable based on the input data, and $D_{KL}(q_{\phi}(z|x)|p(z)),$ is the Kullback-Leibler (KL) divergence this posterior distribution and the prior distribution of the latent variable. Optimizing the loss (Equation 23) involves adjusting generative model ($\theta$) and encoder ($\phi$) parameters [46]. As illustrated in Figure 4, the VAE ’s encoder compresses the channel gain matrix into a lower-dimensional latent code, and the decoder reconstructs it, minimizing reconstruction loss during training.
@@
Using VAEs for resource allocation offers advantages like representing complex distributions, capturing non-linear relationships, and providing a compressed latent space useful for other tasks. They offer a versatile solution to inherently non-convex and challenging resource allocation problems.
The dataset used to train the SS-VAE model consists of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples. These samples are represented by pairs of channel matrices and their corresponding power allocation vectors, denoted as $H_{i}$ and $\Gamma_{i}$ for labeled samples (where $i$ ranges from 1 to $Q_{L}$), and $H_{j}$ for unlabeled samples (where $j$ ranges from 1 to $Q_{U}$). The labeled dataset results from the ESA applied to the same training dataset. Specifically, we have used power transmission and allocation decisions ($p,\alpha,\beta$) obtained by ESA as labels for training our SS-VAE model. The supervised learning component of our algorithm uses the ADAM [47] optimizer and an Mean Squared Error (MSE) loss function for the $Q_{L}$ labeled samples, defined as follows:
@@
The encoder part of the VAE is used to estimate the transmission power and predict the allocation indicators ($\alpha_{u_{s}}^{b}$ and $\beta_{u_{s},m}^{b}$) from the flattened $UB$ dimensional input channel gain matrix.
5.3.2 Contrastive Loss
Contrastive loss, common in DL for representation learning, aims to map similar data points closer and dissimilar ones further apart in the representation space [13]. Mathematically, it can be expressed as:
@@ 5.3.2 Contrastive Loss
where $N$ is the number of data points, $\mathbf{x}_{i}$ is the $i^{th}$ data point, $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ are similar and dissimilar data points, $y_{i}$ is a binary label indicating whether $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$ are similar or dissimilar, $d(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$ is the distance function between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, and $m$ is a margin that separates the similar and dissimilar data points.
To apply contrastive loss in our resource allocation problem, we generate similar and dissimilar channel gain matrices using a channel statistical model that captures the communication environment. Introducing randomness via a Rayleigh fading model produces matrices with varying magnitude and phase coefficients. Specifically, we create a complex Gaussian random matrix with zero mean, where the variance depends on transmitter-receiver distance and accounts for interference as noise [48, 49]. Adjusting this variance controls channel correlation, creating similar or dissimilar matrices. With these, we train our model using the contrastive loss function formulated as follows:
@@
where $\theta$ is the set of parameters in the VAE model, $\tau$ is a temperature hyperparameter, and $L(\theta)$ is the contrastive loss function. $\underline{H_{i}}$ and $\bar{H_{i}}$ are matrices similar to the main channel gain matrix $H_{i}$, used for training the SS-VAE model [50]. Minimizing this loss finds $\theta$ that enables the model to distinguish similar and dissimilar channel gain matrices, thereby addressing the resource allocation problem. The connection between the VAE and the contrastive loss is key for optimal performance. The first step uses supervised learning to learn a good representation of the problem, and the next step refines these representations, improving the model’s generalization to new, unseen scenarios.
5.3.3 Training the SS-VAE model
The SS-VAE training consists of two parts: supervised learning (i.e., training the VAE using its encoder for prediction) and unsupervised learning (i.e., utilizing contrastive loss to enhance generalization and robustness). Key hyperparameters include the latent space dimension (balancing expressiveness and overfitting) [51], learning rate, batch size, and the number of epochs. $L1$ and $L2$ regularization mitigates overfitting [52]. The model uses PyTorch [53], with a 20% validation/test split. Tuned hyperparameters are in Table V.
5.3.4 Performance Metrics
@@ 6 SIMULATION RESULTS
6.1 Simulation Configuration
This section presents the numerical results for the resource allocation problem to evaluate the performance of the SS-VAE model compared to ESA and DQN. In our training and testing scenarios, each RU in the O-RAN deployment is equipped with 25 PRBs and operates with a single antenna with 10 UEs for each service. We consider a network with two slices/services, eMBB and URLLC services. These UEs are randomly distributed within a single-cell environment, mostly 10 to 500 meters from the RU. The initial values and parameters used to run and simulate SS-VAE and the benchmark algorithms (ESA and DQN) are provided in Table V, Table V, and Table V.
For traffic generation, URLLC packets follow a constant bit rate (CBR) model with small fixed payloads of 32 bytes transmitted every 1 ms, resulting in a data rate of approximately 256 kbps per URLLC user. In contrast, eMBB traffic is modeled using a Poisson arrival process with packets sized at 1000 bytes to emulate realistic bursty user behavior such as web browsing and file downloads.
For the channel modeling, we adopt the standard 3GPP-compliant formulation where the channel coefficient between user $u_{s}$ and RU $b$ on PRB $m$ is given by $\mathbf{h}_{u_{s},m}^{b}=\sqrt{\beta_{u_{s},m}^{b}},\mathbf{g}_{u_{s},m}^{b}$. The large-scale fading component $\beta_{u_{s},m}^{b}$ accounts for both path loss and shadowing and is computed as $\beta_{u_{s},m}^{b}=10^{-(\text{PL}_{\text{dB}}+S)/10}$, where the path loss in dB is modeled using the 3GPP Urban Macro (UMa) path loss model which is formulated as $\text{PL}_{\text{dB}}=13.54+39.08\log_{10}(d)+20\log_{10}(f_{c})$, where $d$ is the 2D distance between the UE and RU in meters, and $f_{c}$ is the carrier frequency in GHz. For $f_{c}=2$ GHz, this simplifies to $\text{PL}_{\text{dB}}=13.54+39.08\log_{10}(d)+6.02.$ Shadow fading $S$ follows a log-normal distribution $\mathcal{N}(0,8^{2})$.
The small-scale fading component $\mathbf{g}_{u_{s},m}^{b}$ is modeled as a Rayleigh fading coefficient, i.e., $\mathbf{g}_{u_{s},m}^{b}\sim\mathcal{CN}(0,1)$, to represent non-line-of-sight propagation conditions.
TABLE III: Resource Allocation Simulation Parameters
@@ 6.1 Simulation Configuration
Maximum fronthaul capacity 46 bps/Hz
Maximum data rate for eMBB 10 bps/Hz
Maximum data rate for URLLC 2 bps/Hz
@@
TABLE IV: DQN Training Parameters
@@
$\gamma$, $\epsilon_{\text{final value}}$, $\epsilon_{\text{decay factor}}$ 0.9, 0.1, 0.9995
@@
TABLE V: SS-VAE Model Training Parameters